A jet aircraft with mass of 582,731 kilograms is flying straight upward. At an instant the plane's engine is producing an upward thrust (force) of 2,224,800 Newtons and air resistance is exerting a force of 90,000 Newtons downward. What is the acceleration of the plane in m/s2 at that instant? Note: In this problem, we define "up" as positive.
Answer #1

Using Force balance on aircraft in vertical direction:

F_net = Ft - Fr - W

Ft = Thrust force by engine in upward direction = 2224800 N

Fr = Resistive force by air in downward direction = 90000 N

W = Weight of aircraft in downward direction = m*g = 582731*9.8

F_net = net force on aircraft from Newton's 2nd law = m*a

So,

m*a = Ft - Fr - W

a = (Ft - Fr - W)/m

a = (2224800 - 90000 - 582731*9.8)/582731

a = -6.1365 m/s^2 = acceleration of aircraft at this instant

(-ve sign means aircraft is decelerating)

Let me know if you've any query.